# Multicopy Reinforcement Learning Agents

### v2

@@ Abstract
This paper examines a novel type of multi-agent problem, in which an agent makes multiple identical copies of itself in order to achieve a single agent task better or more efficiently. This strategy improves performance if the environment is noisy and the task is sometimes unachievable by a single agent copy. We propose a learning algorithm for this multicopy problem which takes advantage of the structure of the value function to efficiently learn how to balance the advantages and costs of adding additional copies.
[…]
2695
1 Introduction
@@
In most multi-agent domains, agents either cooperate or compete to achieve their goals [6, 4, 15]. This paper seeks to address a different type of multi-agent scenario, one in which agents duplicate themselves to execute multiple copies of the same policy to combat noise and risk in the problem domain. One such domain is packet forwarding in mobile wireless networks, where duplicate packets have been shown to improve reliability and reduce delay [10] as well as other domains, like robot search and rescue, where duplicate agents may improve performance.
We assume a task that, in a noiseless and no risk environment, could be completed by a single agent following an optimal policy. However, when noise increases to a sufficiently high level a single agent may fail to complete the task. In this case, an agent that can duplicate itself and make several attempts concurrently to achieve the goal may be more successful, or able to complete the task more quickly.
We will assume for now that (1) the only coordination between agents occurs when new copies are made at the start state, and that (2) after that point each agent copy operates completely independently of the others. We will also assume that (3) aside from the points at which copies are made the agents all follow identical policies. These three assumptions will give a particular form to the value function for the agent group, a form which factors easily and has many of the properties of a single agent value function, without requiring consideration of the full cross product of agent states and actions even in the exact tabular case.
2 Background: Reinforcement Learning
@@
Reinforcement Learning (RL) assumes an agent existing in an environment which can be represented by a Markov Decision Process (consisting of states $S$, actions $A$, and transition probabilities between states) [12]. The agent seeks to maximize a reward signal over sequential, extended time interaction with the environment.
@@
$G_{t}=r_{t}+\gamma\cdot r_{t+1}+\dots+\gamma^{T-t}r_{T}$
@@
There are two types of algorithm that can be used to learn how to behave in this type of environment: bootstrapping algorithms, which update the $Q$ value of each state and action based on the current estimate of the value of the next state(s) (e.g. Q-learning [14] and SARSA [7, 11] and non-bootstrapping algorithms, which work directly from samples of $G_{t}$ (e.g. Markov Chain Monte Carlo [2, 9]).
3 Background: Cooperative Multiagent RL
Multiagent Reinforcement learning often starts by considering an MDP that incorporates the states and actions of each agent in a set of agents. In the case of cooperating agents (the case closest to the problem we consider here), the agents are considered to be optimizing a joint reward function in this combined multi-agent MDP. Within this MDP, the number of states and actions grows exponentially with the number of agents. Work in multi-agent RL typically uses function approximation to simplify the large joint state space, or uses a simplified view of the joint state that is only partially observed by each agent or small group of agents, as in e.g. [5, 8].
In this work, we take a different approach. While in the future we will add function approximation, we begin by examining whether we need to consider the joint state and action space at all, in order to start with a problem formulation that is as simple as possible. We conclude that we do not need the joint state space, though we do need to consider the joint actions when making choices between them.
4 Basic Multiagent Algorithm Baseline
The problem we outline in this paper does not translate immediately to a standard tabular multiagent solution. In order to test against a realistic baseline, we have designed a multiagent solution with a basic communication protocol and a reward function that reflects the cooperative and competitive aspects of the task. As a multiagent problem, the task has the following properties: (1) Agents may only communicate when they split into multiple copies, in the start state. (2) Only the best agent to achieve the goal gets rewarded for doing so. (3) Costs are summed across all agents. This means that while the overall task is cooperative, individual agents compete to best achieve the goal and receive the reward for doing so.
To reflect these constraints, we implemented independent agents with communication and coordination in the start state. All agents receive negative cost rewards as they go, but only the best agent receives positive optimization reward at the end of the task.
Each agent remembers the joint action chosen at the start state. If $n$ agents choose the action set $\{a_{0},a_{1},\dots,a_{n}\}$ in the start state, then at every state $s$ in the episode we add this information to the state, forming the augmented state $(s,(a_{1},a_{2},\dots a_{n}))$. This provides enough information to estimate the agent’s probability of success in the race for the optimization reward in the context of the other agents, but increases the size of the state space. Once the state space is augmented in this way, each agent runs q-learning with a shared value function between all agents.
In the start state, when calculating the value of a candidate set of actions $(a_{1},a_{2},\dots a_{n})$, the values for each individual action in the context of the joint action are summed. The action combination with the highest value is chosen. From that point forward, the agents act independently, even if they are on similar paths.
5 Variance and Action Combinations
It may seem natural to order actions by their expected value when deciding which paths to send agents down. However, when combining multiple agents to find the highest performing one, actions with high variance may combine better than agents with low variance.
Consider a simple Multiarmed Bandit [12] problem with two arms, arm 1: normal distribution with $\mu=10,\sigma=1$ and arm 2: normal distribution with $\mu=5,\sigma=30$. Mean sample values for the maximum reward returned by combinations of up to 2 arms after 4000 sample, are highest when the noisiest arm is used twice, shown as follows:
Action 1 2 (1,1) (1,2) (2,2)
Est. Val. 10.0 5.22 10.56 19.70 22.11
@@ 5 Variance and Action Combinations
Action (2, 2) is the best 2-action combination when the maximum of the two arms is used as the overall return. Due to it higher variance, it is more likely to generate combinations of outcomes that have at least one high return.
6 Multicopy Actions
@@
At certain states the agent may make up to $n$ copies of itself. In state $s$, considering multi-actions that include up to $n$ actions, the set of multi-action combinations available is:
@@
(a) Three Bridges Gridworld: quick, risky bridge (A); slow, safe bridge (B); medium bridge (C).
(b) Broken Bridges Gridworld with one broken bridge (A). The broken bridge is randomly chosen per episode.
Figure 1: Three-bridge gridworlds. Terminal states are marked with double squares. All other states have 4 actions: North, South, East, West. Postive terminal state rewards are optimization rewards, negative are costs. White squares have a per-step cost that is set for each experiment.
@@
Optimization rewards are only counted on the “best" path found. Therefore at each timestep they are maximized over the returns from the next timestep. Let $R_{o}(t,i)$ be the optimization reward at time $t$ for agent $i$, and $G_{o}(t,i)$ be the corresponding optimization return. When calculating overall optimization reward for the set of agent copies, we only count the best agent’s return:
@@
$Q(s,$ $m)=E\left[G(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$
$=E\left[G_{c}(t,i)+G_{o}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$
@@
$Q(s,m)=~{}$ $E\left[G_{c}(t,i)+G_{o}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$
@@
$Q_{c}(s,m)$ $=E\left[G_{c}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$
$=E\left[\sum_{j\in ID_{t}(i)}G_{c}(t,j)\mid S_{t,i}=s,\in M_{t,i}=m\right]$
@@
$=\sum_{a\in m}E\left[G_{c}(t,j)\mid S_{t,i}=s,A_{t,j}=a\right]$
$=\sum_{a\in m}Q_{c}(s,a)$
Therefore, for $Q_{c}$ we can use any learning technique, including bootstrapping methods like Q-learning [14, 12] or SARSA [7, 12] to learn the Q-function for individual actions, and sum for multiactions.
@@
$Q_{o}(s,m)$ $=E\left[\max_{j\in ID_{t}(i)}G_{o}(t,j)\mid S_{t,i}=s,M_{t,j}=m\right]$
Here we must learn $Q_{o}$ for each combination of actions we consider. However, within the maximization step each $G_{o}$ may be calculated individually, allowing us to avoid a value function defined over the cross-product of states. If we are estimating $G_{o}$ over multiple times steps this does limit us to techniques which do not use bootstrapping like Markov Chain Monte Carlo [12].
Policy Improvement
When calculating the best action, the agent should optimize the total value:
$\pi$ $(s)=$
$\operatorname*{arg\,max}_{m}E\left[G_{c}(t,i)+G_{o}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right].$
@@
7 Multicopy Algorithm
Multicopy agent
The multicopy agent combines a cost agent ($Q_{c}$), and an optimization agent ($Q_{o}$), and sums their values to calculate the joint multicopy policy $\pi$. This joint policy is used by both cost and optimization agents.
Cost and Optimization Agents
The cost agent estimates $Q_{c}$ for each state $s$ and individual action $a$ using Q-learning. To calculate the value of a multiaction $m$, it sums the values of its component actions. The optimization agent estimates $Q_{o}$ for each state $s$ and multiaction $m$, using Every Visit MCMC.
8 Gridworld Domains
@@
Gridworld Environment. To model tasks with varying risk, delay and cost in a simple domain, we use a set of gridworld bridges of varying length and width, shown in Figure 1(a). This is an expanded version of a single bridge gridworld from the Pacman Projects [3]. Falling off a bridge represents task failure, and successfully making it across any bridge represents task success.
The three bridges present different trade-offs between speed and risk. Bridge (A) is short but narrow, useful when noise is low but risky when noise is high. Bridge (B) is long and wide: safer, but slower. Bridge (C) is a compromise.
Start state
In the start state (not shown), the agent may choose any single bridge or a combination of bridges, up to a maximum number of actions (set to 1, 2 or 3 per experiment). The actions in the start state have no noise, transitioning to the states labeled "A", "B" or "C" in Figure 1(a). An agent copy cannot transition back to the start state once it has chosen a bridge.
Noise
The gridworld includes an adjustable noise parameter [3]. With noise setting $\beta$, directional actions (N, S, E, W) on the bridges go in the intended direction with probability $1-\beta$, and go in the two orthogonal directions with probability $\beta/2$ each.
Optimization and Costs
The optimization reward, shown as positive reward in Figure 1(a), rewards successful crossings. The cost reward consists of a small negative reward on each step in the white bridge states, and a larger negative reward (shown) for falling off the bridge. The cost in bridge states varies per experiment.
(a) Basic learning Comparison: Noise 0.1
@@
(c) Long term convergence: Noise 0.1
(d) Basic learning Comparison: Noise 0.3
@@
(f) Long term convergence: Noise 0.3
Figure 2: Three Bridges Learning results for basic non-multicopy vs multicopy algorithms on the 3 bridges gridworld. Cost per step is -2 in all plots. Shaded lines are averaged over 50 trial runs, dark lines are rolling 30-episode averages of those values.
Broken Bridges
In the version of the three bridge gridworld shown in Figure 1(b), at the start of each episode one randomly chosen bridge is broken midway along (Bridge A in the diagram). All agents crossing the broken bridge during the episode will fail.
Exploration
On-policy techniques like MCMC can be overly cautious when noise is high. We therefore use a boltzmann distribution with action advantages and decreasing temperature over time, starting at 100, for exploration.
Algorithm Parameters
The learning rates for MCMC and Q-learning typically differ, with lower learning rates for MCMC. In all algorithms we use 0.05 for MCMC and 0.2 for Q-learning, linearly decreasing to 0 through the training episodes. We used a discount of 0.9 for all experiments.
9 Experiments
The experiments in this section aim to determine if the algorithm can find the right number of agent copies and combination of actions. First, we compare our algorithm from Section 7 to the basic algorithm with joint actions from Section 4 in the Three Bridges Gridworld, with various maximum action number settings. Then we compare the number and type of actions chosen for various cost and noise settings, to confirm that the algorithm is able to adjust the multiactions chosen according to the domain. Next we examine the difference in policy when we allow duplicate actions (more than one agent on the same bridge), and finally, some more nuanced experiments with duplicate actions.
Training and Testing
Each agent was trained for 7500 episodes and tested for 500 episodes, except where noted. During the testing phase, exploration and learning are turned off completely and only the optimal multiaction is chosen.
(a) Return. Per Episode return during testing phase for various noise and cost settings, after 7500 episodes of training. Final graph shows proportion improvement from the Joint Action return to the Multicopy return.

### v3

@@ Abstract.
Inspired by the problem of optimal packet duplication in Mobile Wireless Networks, this paper examines a specific type of cooperative multi-agent problem in which an agent makes multiple identical copies of itself in order to achieve a noisy and difficult single agent task more reliably or efficiently. The agent must balance the cost of sending more copies with the improvement in speed or success rate generated by the extra copies. Due to the space of possible joint reward functions, this specific case has some differences from the more general cooperative multi-agent problem. We propose a learning algorithm for this multicopy problem which takes advantage of the structure of the value function to efficiently learn how to balance the costs and benefits of adding additional copies.
Key words and phrases:
Multiagent Reinforcement Learning, Cooperative Agents, Wireless Networks
1. Introduction
This paper introduces a new type of multiagent problem, in which a single agent duplicates itself to achieve a single goal with greater speed, quality or reliability. While this problem can be cast as a specific case of cooperative agents (for algorithms in which the number of agents can be optimized), the problem has properties that lend themselves to solutions that are tuned for this scenario.
@@ 1. Introduction
In most multi-agent domains, agents either cooperate or compete to achieve their goals Albrecht et al. (2024); Gronauer and Diepold (2022); Wong et al. (2023). This paper seeks to address a different type of multi-agent scenario, one in which agents duplicate themselves to combat noise and risk in the problem domain. One such domain is packet forwarding in mobile wireless networks, where duplicate packets have been shown to improve reliability and reduce delay Spyropoulos et al. (2008). Other domains include robot search and rescue, where duplicate agents may improve performance.
We assume that, in a noiseless and no risk environment, the task at hand could be completed by a single agent following an optimal policy. We further assume that if additional agents complete the task, this does not improve the outcome. However, actions are so noisy that even an agent behaving optimally has a nontrivial chance of failure. In this case, an agent may duplicate itself and make several attempts concurrently to achieve the goal. This agent must balance two factors: the cost of adding more copies vs the increase in performance due to replication. In this case, the reward function factors into a cost portion, which is summed over all agents, and an optimization portion, which is taken from only the highest performing agent. This restricts the form of the reward function of what is otherwise a cooperative multiagent setting.
We also assume that communication between agents is prohibitively expensive, except at the point where agent copies are made. When copies are made the agent(s) may jointly decide how many copies to make and what combination of actions to take: whether to take the same action with multiple copies or a range of actions to diversify the copies. After this point, we assume that noise in the domain will suffice to produce a diverse set of outcomes for different copies.
2. Application to Wireless Networks
In Mobile Wireless Networks, particularly sparse networks, there are cases in which not all packets make it to their destination within a reasonable time frame Jain et al. (2004). A mobile wireless network can be viewed as a dynamic graph over a time. In a dynamic network, contiguous paths from a source to a destination may exist instantaneously or only across time. In this situation it can be useful to send more than one copy of a packet Spyropoulos et al. (2008), so that at least one copy gets to the destination in a reasonable time frame. We have designed our test bed to duplicate this scenario in some respects.
As in Manfredi et al. ([n.d.]), we treat each packet as an agent moving through an environment made up of devices. There are some properties of the Wireless Network we have constructed our gridworld to match: messages/agents must travel from a source to a goal point; only one copy of the agent must reach the goal for success and success among the copies varies due to noise in the environment.
To construct a simple gridworld in which to test our algorithms, we made some simplifying assumptions: we use tabular states and actions rather than feature based RL with machine learning; in the mobile network, the available actions will change as device neighbors change, whereas in our gridworld the action set in each state is fixed; and packets affect each other indirectly through congestion in the network: actual network data will include features that reflect current network congestion levels, but no such effect is modeled in our test scenario.
3. Reinforcement Learning
@@
Reinforcement Learning (RL) assumes an agent existing in an environment which can be represented by a Markov Decision Process (consisting of states $S$, actions $A$, and transition probabilities between states) Sutton and Barto (2018). The agent seeks to maximize a reward signal over sequential, extended time interaction with the environment.
@@
$G_{t}=r_{t}+\gamma\cdot r_{t+1}+\dots+\gamma^{T-t}r_{T}$ (1)
@@
4. Best Actions vs Best Multiaction
Combining the cost reward for our agent copies is a straightforward sum. The optimization reward, on the other hand, is maximized over all agent copies, resulting in more interesting behavior. When splitting into multiple copies, the agent may choose a set of immediate actions for the copies to take. Always including the action with the highest individual expected value may seem best. However, a combination of actions with high expected values but low variances can lead to worse outcomes than a combination with lower expected values but higher variance.
Consider a simple Multiarmed Bandit Sutton and Barto (2018) problem with two arms, arm Stable (S): normal distribution with $\mu=10,\sigma=1$ and arm Noisy (N): normal distribution with $\mu=5,\sigma=30$. To get the expected return of action sets like (S, N), we can generate samples from both distributions and take the max of the samples. Mean optimization reward returned by combinations of up to 2 arms after 10,000 samples are shown in Table 1, rounded to whole numbers.
Table 1. Max of stable (S) and noisy (N) normal distributions.
Action S N (S,S) (S,N) (N,N)
Est. Val. 10 5 11 20 22
@@ 4. Best Actions vs Best Multiaction
Action (N, N) is the best 2-action combination, despite action N having a lower expected value. Due to higher variance, two copies of arm N are more likely to generate at least one high return.
In Table 2 we can see the same experiment for a bandit in which in which a mix of two different distributions is best: the two arms are a constant (C) with value 100, and exponential (E) $exp(1/70)$.
Table 2. Max of constant (C) and exponential (E) distributions.
Action C E (C,C) (C,E) (E,E)
Est. val. 100 70 100 116 104
Here (C, E), a stable distribution combined with a noisy distribution provides the best expected maximum.
5. Cooperative Multiagent RL
Multiagent Reinforcement learning often starts by considering an idealized MDP that uses the cross product of the states and actions of each agent in a set of agents. In the case of cooperating agents, the agents optimize a joint reward function in this multi-agent MDP Albrecht et al. (2024). Work in multi-agent RL typically uses function approximation to simplify the large joint state space, or uses a view of the joint state that is only partially observed by each agent or small group of agents, as in e.g. Khorasgani et al. (2021); Schroeder de Witt et al. (2019). We instead begin by examining whether we need to consider the joint state and action space at all. Once copies have been created, they independently work to achieve the goal to the best of their ability. As we will see in Section 7, there is no coordination needed outside joint actions in states where copying occurs.
Figure 1. Shadowed Equilibrium Example.
Shadowed Equilibria. Multicopy RL shares the problem of shadowed equilibria Oroojlooy and Hajinezhad (2023); Fulda and Ventura (2007) with more general cooperative RL scenarios, however, the problem occurs differently. Consider the example MDP shown in Figure 1. In a general cooperative setting we can construct a shadowed equilibrium as follows: (1) place two agents at states $S_{1}$ and $S_{2}$; (2) set the reward on the action combinations to, e.g.: $(a_{1}^{*},a_{2}^{*})=10$, $(a_{1},a_{2})=5$, $(a_{1}^{*},a_{2})=1$, $(a_{1},a_{2}^{*})=1$; and (3) have many $a_{1}$ and $a_{2}$ distractor actions available. In algorithms where the two agents act independently, during random exploration, $a_{1}$ will have better estimated value than $a_{1}^{*}$ because both are more likely to be paired with one of the $a_{2}$ distractors. Similarly, $a_{2}$ will have better estimated value than $a_{2}^{*}$. Once the action combination $(a_{1},a_{2})$ has been learned, the agent is stuck: trying $a_{1}^{*}$ or $a_{2}^{*}$ looks even worse than during random exploration.
To construct this scenario in our optimization reward case, we cannot simply set the rewards: they are set indirectly through the maximization of of individual action reward distributions. However, we can indirectly construct reward distributions for the 4 actions by modifying the example from Table 2. Higher mean distributions go to the starred actions: $a_{1}^{*}=110$, $a_{2}^{*}=10+exp(1/70)$. Lower distributions go to the distractor actions: $a_{1}=exp(1/70)$, $a_{2}=100$.
Table 3. High and low mean constant and exponential distributions.
Action $(a_{1}^{*},a_{2})$ $(a_{1}^{*},a_{2}^{*})$ $(a_{1},a_{2})$ $(a_{1},a_{2}^{*})$
Est. Val. 110 127 117 111
When we combine these distributions and take the expected maximum, constants combined with exponentials give the highest values, as shown in Table 3. This satisfies the requirements for a shadowed equilibrium to occur when there are enough distractor actions.
6. Baseline Multiagent Algorithm
As a baseline to compare to our algorithm (see Section 7), we implemented a multiagent solution with a basic commuication protocol and reward function, referred to as the Joint Action Multiagent algorithm. As a standard multiagent problem, the task has the following properties: (1) Agents may only communicate when they split into multiple copies. (2) Only the best agent to achieve the goal gets the reward for doing so. (3) All agents accumulate cost rewards. This means that while the overall task is cooperative, individual agents compete to best achieve the goal.
While this reward function accurately reflects the task, the reward received by an agent depends on the other agents: if agent copy 2 does better at the task, then agent copy 1 receives no reward for success. Therefore, the number of copies made and the actions they took in the splitting state affect the value function for each agent. Each agent must remember the joint action chosen at the start state, or operate in a POMDP. If $n$ agents choose the action set $\{a_{0},a_{1},\dots,a_{n}\}$ in the start state, then at every state $s$ in the episode we add this information to the state, forming the augmented state $(s,(a_{1},a_{2},\dots a_{n}))$. Once the state space is augmented, each agent runs q-learning with a shared value function between all agents. When calculating the value of a candidate set of actions $(a_{1},a_{2},\dots a_{n})$, the values for each individual action in the context of the joint action are summed. The action combination with the highest value is chosen. From that point forward, the agents act independently, even if they are on similar paths.
7. Multicopy Actions
Ideally, we’d like each agent to attempt the task to the best of its ability, despite the presence of other copies. We would therefore like to consistently reward each copy for learning to achieve the goal. However, when calculating the value of a joint action, we should only include the optimization reward received by the best copy. Our algorithm does so, and allows us to treat the domain as an MDP without remembering the joint action chosen.
@@ 7. Multicopy Actions
At certain states the agent may make up to $n$ copies of itself. In state $s$, considering multi-actions that include up to $n$ actions, the set of multi-action combinations available is the set of multisets of cardinality less than $n$ that can be formed from the set of actions:
@@
Optimization rewards are only counted on the “best” path found. Therefore at each timestep they are maximized over the returns from the next timestep. Let $R_{o}(t,i)$ be the optimization reward at time $t$ for agent $i$, and $G_{o}(t,i)$ be the corresponding optimization return. When calculating overall optimization reward for the set of agent copies, we only count the best agent’s return:
@@
$Q(s,m)=\leavevmode\nobreak$ $E\left[G(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$ (10)
[…]
@@
$Q(s,m)=\leavevmode\nobreak$ $E\left[G_{c}(t,i)+G_{o}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$ (12)
@@
$Q_{c}(s,m)=\leavevmode\nobreak$ $E\left[G_{c}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right]$ (16)
[…]
@@
[…]
Therefore, for $Q_{c}$ we can use any learning technique, including bootstrapping methods like Q-learning Watkins and Dayan (1992); Sutton and Barto (2018) or SARSA Rummery and Niranjan (1994); Sutton and Barto (2018) to learn the Q-function for individual actions, and sum for multiactions.
@@
$Q_{o}(s,m)=$ $\leavevmode\nobreak\ E\left[\max_{j\in ID_{t}(i)}G_{o}(t,j)\mid S_{t,i}=s,M_{t,j}=m\right]$ (20)
Here we must learn $Q_{o}$ for each combination of actions we consider. However, within the maximization step each $G_{o}$ may be calculated individually, allowing us to avoid a value function defined over the cross-product of states. If we are estimating $G_{o}$ over multiple times steps this does limit us to techniques which can be used without bootstrapping like Markov Chain Monte Carlo Sutton and Barto (2018).
Policy Improvement. When calculating the best action, the agent should optimize the total value:
$\pi(s)=$ $\operatorname*{arg\,max}_{m}E\left[G_{c}(t,i)+G_{o}(t,i)\mid S_{t,i}=s,M_{t,i}=m\right].$ (21)
@@
Multicopy agent. The multicopy agent combines a cost agent ($Q_{c}$), and an optimization agent ($Q_{o}$), and sums their values to calculate the joint multicopy policy $\pi$.
Cost and Optimization Agents. The cost agent estimates $Q_{c}$ for each state $s$ and individual action $a$ using Q-learning. To calculate the value of a multiaction $m$, these are summed. The optimization agent estimates $Q_{o}$ for each state $s$ and multiaction $m$, using Every Visit Markov Chain Monte Carlo Sutton and Barto (2018).
(a) Three Bridges: quick, risky bridge (A); slow, safe bridge (B); medium bridge (C).
(b) Bridges Gridworld with one broken bridge (A). The broken bridge is randomly chosen per episode.
Figure 2. Three-bridge gridworlds. Terminal states are marked with double squares. All other states have 4 actions: North, South, East, West. Postive terminal state rewards are optimization rewards, negative are costs. White squares have a per-step cost that is set for each experiment.
8. Gridworld Domains
@@
Gridworld Environment. To model tasks with varying risk, delay and cost in a simple domain, we use a set of gridworld bridges of varying length and width, shown in Figure 2(a). This is an expanded version of a single bridge gridworld from the Pacman Projects DeNero and Klein (2010). Falling off a bridge represents task failure, and successfully making it across any bridge represents task success.
The three bridges present different trade-offs between speed and risk. Bridge (A) is short but narrow, risky when noise is high. Bridge (B) is long and wide: safer, but slower. Bridge (C) is a compromise.
Start state. In the start state (not shown), the agent may choose any single bridge or a combination of bridges, up to a maximum number of actions (set to 1, 2 or 3 per experiment). The actions in the start state have no noise, transitioning to the states labeled ”A”, ”B” or ”C” in Figure 2(a). An agent copy cannot transition back to the start state once it has chosen a bridge.
Noise. The gridworld includes an adjustable noise parameter DeNero and Klein (2010). With noise setting $\beta$, directional actions (N, S, E, W) on the bridges go in the intended direction with probability $1-\beta$, and go in the two orthogonal directions with probability $\beta/2$ each.
(a) Basic learning: Noise 0.1
@@
(c) Long term: Noise 0.1
(d) Basic learning: Noise 0.3
@@
(f) Long term: Noise 0.3
Figure 3. Three Bridges: Learning results for basic non-multicopy vs multicopy algorithms on the 3 bridges gridworld. Cost per step is -2 in all plots. Shaded lines are averaged over 50 trial runs, dark lines are rolling 30-episode averages of those values.
Optimization and Cost. The optimization reward, shown as positive reward in Figure 2(a), rewards successful crossings. The cost reward consists of a small negative reward on each step in the white bridge states, and a larger negative reward (shown) for falling off the bridge. The cost in bridge states varies per experiment.
Broken Bridges. In the version of the three bridge gridworld shown in Figure 2(b), at the start of each episode one randomly chosen bridge is broken midway along (Bridge A in the diagram). All agents crossing the broken bridge during the episode will fail.
9. Experiments
The experiments in this section aim to determine if the algorithm can find the right number of agent copies and combination of actions. First, we compare our algorithm from Section 7 to the basic algorithm with joint actions from Section 6 in the Three Bridges Gridworld, with various maximum action settings. Then we compare the number and type of actions chosen for various cost and noise settings, to confirm that the algorithm is able to adjust the multiactions chosen according to the domain. Next we examine the difference in policy when we allow duplicate actions (more than one agent on the same bridge), and finally, some more nuanced experiments with duplicate actions.
Exploration. On-policy techniques like MCMC can be overly cautious when noise is high. We therefore use a boltzmann distribution with action advantages and linearly decreasing temperature over time, starting at 100 and ending at 1, for exploration. Temperature decreases at the end of each episode.
Algorithm Parameters. The learning rates for MCMC and Q-learning typically differ, with lower learning rates for MCMC. In all algorithms we use 0.05 for MCMC and 0.2 for Q-learning, linearly decreasing to 0 through the training episodes. We used a discount of 0.9 for all experiments.
Training and Testing. Each agent was trained for 7500 episodes and tested for 500 episodes, except where noted. During the testing phase, exploration and learning are turned off completely and only the optimal multiaction is chosen.
9.1. Learning Curves
In Figure 3(b -c) and (e-f), we compare Multicopy and Joint Action algorithms with various maximum action settings. Figures (b) and (e) use the 7500 episode training time that we will use throughout the remainder of the paper, while Figures (c) and (f) use a longer training time of 50,000 episodes. The Multicopy algorithm does almost as well in 7500 episodes as it does in 50,000, even with high noise levels. The Joint Action algorithm, however, due to its fractured state space, does not learn as well when training time is short, particularly in (e), where noise is high.
(a) Return. Per Episode return during testing phase for various noise and cost settings. Final graph shows proportion improvement from the Joint Action return to the Multicopy return.